OpenStudy (anonymous):

Verify that f and g are inverse functions mathematically : f(x)=(3-x)/ x g(x)=(3)/x+1 solve: f(g(x))

OpenStudy (helder_edwin):

u have \[ \large f(x)=\frac{3-x}{x}\qquad g(x)=\frac{3}{x+1} \] right?

OpenStudy (anonymous):

yes.

OpenStudy (helder_edwin):

to verify they r inverses of each other just do this \[ \large (f\circ g)(x)=f(g(x)) \] and \[ \large (g\circ f)(x)=g(f(x)) \]

OpenStudy (anonymous):

f(g(x))=f((3)/x+1)=(3-(3)/x+1)/ ( (3)/x+1)= x+1-1 =x

OpenStudy (helder_edwin):

u should get \(x\) when finished

OpenStudy (anonymous):

Yea thats what i did but for f(g(x)) i somehow got 1so im going to redo it but i did get X for g(f(x))

OpenStudy (helder_edwin):

let's do it then

OpenStudy (helder_edwin):

\[ \large (f\circ g)(x)=f(g(x))=f\left(\frac{3}{x+1}\right)= \] \[ \large =\frac{3-\frac{3}{x+1}}{\frac{3}{x+1}}= \frac{\frac{3(x+1)-3}{x+1}}{\frac{3}{x+1}} \]

OpenStudy (helder_edwin):

ok so far?

OpenStudy (anonymous):

hmm how did you get the numerator to equal that? I thought you would have to rationalize 3 to have a denominator of x+1

OpenStudy (helder_edwin):

u do understand the substitution, though. right?

OpenStudy (anonymous):

Yes I got that.

OpenStudy (helder_edwin):

now in the numerator i just added the fractions: \[ \large a+\frac{b}{c}=\frac{ac+b}{c} \]

OpenStudy (anonymous):

so you multiplyed -1 * x+1 then added 3?

OpenStudy (helder_edwin):

i don't understand your question.

OpenStudy (helder_edwin):

i multiplied 3 times x+1 and -1 times 3

OpenStudy (helder_edwin):

well after simplifying \[ \large =\frac{3x+3-3}{3}=\frac{3x}{3}=x \]

OpenStudy (anonymous):

oh ok I see sorry i had put the input for x in parenthesis which made it look like 3-1(3/x+1) and ok yea i got that now thanks!

OpenStudy (helder_edwin):

u r welcome